a rectangular piece of cardboard has an area of 150 square centimeters. By cutting a square 2 centimeters wide on each side each of the four corners and folding up the sides, an open box body is formed having a volume of 132 cubic centimeters. Find the length and width of the original cardboard

Answer by stanbon(75887)

a rectangular piece of cardboard has an area of 150 square centimeters. By cutting a square 2 centimeters wide on each side each of the four corners and folding up the sides, an open box body is formed having a volume of 132 cubic centimeters. Find the length and width of the original cardboard
-------
Original DATA:
L*W = 150 cm^2
-----
New DATA:
(L-4)(W-4)*2 = 132
-------
(L-4)(W-4) = 66
LW -4L-4W+16 = 66
150 -4L -4W = 50
-------
4(L + W) = 100
L+W = 25
===============================
Substitute for "L" and solve for W::
(25-W)(W) = 150
------
W^2 -25W +150 = 0
---
(W-15)(W-10) = 0
W = 15
L = 10
---------------
Cheers,
Stan H.
------------